3(3x-2) = 5 - 2(2x-3)
------- -------
10 5

This should read 3(3x-2) over 10 is equal to 5 minus 2(2x-3) over 5. I would appreciate the view of the workings as I need to see where I am going wrong.

Answer by stanbon(75887)

3(3x-2) = 5 - 2(2x-3)
------- -------
10 5
-----------------------
Multiply thru by 10 to get rid of the denominators:
3(3x-2) = 50 - 4(2x-3)
Use the distributive law to get:
9x-6 = 50 -8x+12
Combine the x-terms on one side and the numbers on the other:
17x = 78
Divide both sides by 17 to get:
x=4.588...
Cheers,
Stan H.